if it takes 9 men 15 days to do a job, how do you work out how long job will take, if you add another 3 men to help who are twice as fast as those 9 men?
Answer by stanbon(75887)

if it takes 9 men 15 days to do a job, how do you work out how long job will take, if you add another 3 men to help who are twice as fast as those 9 men?
---
# of men and # of days to do the job are indirectly related.
m = k/d
---
Solve for "k" using "it takes 9 men 15 days to do a job"
9 = k/15
k = 135
----
Equation:
m = 135/d
how long will the job take, if you add another 3 men to help who are twice as fast as those 9 men
---
Those 3 men are equivalent to 6 of the other men.
Adding those 3 to the 9 gives you the equivalent of 9+6 = 15 men
---
Solve
m = 135/d with m = 15
15 = 135/d
d = 135/15
d = 9 days (The job can be done in 9 days)
===================
Cheers,
Stan H.